### v2

@@ 3.3 Evaluation strategy
Out-of-sample evaluation is a cornerstone of time series forecasting, providing a way to test a model’s generalization ability beyond the training period, an essential step given that future data may diverge from historical patterns due to structural breaks, level shifts, or unexpected shocks (Tashman, 2000). Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir & Benítez, 2012). It respects the chronological nature of time series data while enabling repeated assessments across multiple forecast origins. In this framework, the time series is split into a training set and a test set, where the model is trained on the former and evaluated on the latter. Forecasts are generated for a defined horizon $h$, and at each iteration, the forecast origin is shifted forward by a specified step size, typically one, to simulate real-time forecasting. The model is then retrained on the updated training data, using either a fixed-length or expanding window. Evaluation metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
Compared to fixed origin evaluations, rolling origin offers a more nuanced view of a model’s robustness by exposing it to a range of temporal conditions (including seasonal patterns, trends, and potential anomalies) that may not be captured in a single evaluation window (Bergmeir & Benítez, 2012). This iterative approach reduces the risk of overfitting to one specific train-test split, and is especially beneficial in operational contexts like retail, logistics, or finance, where forecasts are regularly updated in response to incoming data. It also aligns well with real-world decision-making processes that require forecasts to adapt over time.
In practice, the design of a rolling origin evaluation depends on several parameters: the size of the test set, the length of the forecast horizon, the step size, and the windowing strategy (fixed vs expanding). The expanding window strategy is particularly suited for short time series, as it maximizes the amount of historical data available at each iteration. This approach is widely used in applied settings (Petropoulos & et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series. The Table 1 outlines the key parameters used in our experiments, including a test set spanning at least one full year, to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency. A step size of one is also relevant in the context of stability evaluation, since multiple forecasts for the same target are required to compute stability metrics.
@@
Furthermore, to investigate the impact of retraining on forecasting stability, we examined various retraining scenarios, or retraining windows, as defined in Zanotti (2025b). Each scenario $r$ indicates the number of new observations after which the model is retrained. These retraining scenarios are tailored to the data frequency, daily for M5 and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles. For instance, in the daily case, $r=7$ reflects weekly updates. Across all scenarios, we trained global models on aligned datasets, updating the model either completely at each retraining step or not at all. We excluded hyperparameter tuning due to its high cost and marginal expected benefit. Following Zanotti (2025b), we treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
@@ 3.4.1 Stability metrics
Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and n1n-1 respectively. The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values. Lower values imply less unstable predictions.
Beyond point forecasts, our study also emphasizes the probabilistic stability of the models under different retraining strategies. Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts. Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying retraining scenarios 222 One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, or to use inherently probabilistic models. These are also valid ways of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022). However, comparing different probabilistic modeling frameworks is out of the scope of our analysis, and it may be an area of future research. . Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic quantile forecasts, we proposed a new measure, the Multi-Quantile Change (MQC), to comprehensively assess the stability of the probabilistic predictions. We defined the Quantile Change and the Multi-Quantile Change as:
@@
where $\mathcal{Q}$ is the set of possible quantile levels $\alpha$. The Quantile Change (QC) metric is a measure of the change in forecasted quantiles, $\hat{q}_{t}^{(\alpha)}$, between two consecutive forecast origins. Given two forecast origins, $n$ and n1n-1, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$. The formula mirrors the Quantile Loss (Pinball Loss) but replaces the actual observations $y_{t}$, with previous quantile forecasts $\hat{q}_{t,n-1}^{(\alpha)}$, thereby shifting the focus from forecast accuracy to forecast stability. The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{q}_{t,n}^{(\alpha)}\approx\hat{q}_{t,n-1}^{(\alpha)}$), then QC will be small. Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of stability in its probabilistic output. By aggregating this across a set of quantile levels $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability, providing a comprehensive metric for probabilistic stability.
@@
To further contextualize the proposed metric, it is important to compare it with existing distributional distance metrics such as the Kullback–Leibler ($KL$) divergence and the Wasserstein distance ($W_{2}$). KL divergence measures the relative entropy between two probability distributions and is particularly sensitive to differences in the tails (Kullback & Leibler, 1951). However, it requires access to full predictive densities, which can limit its use in forecasting applications where outputs are often expressed only as quantiles. Wasserstein distance ($W_{2}$), by contrast, quantifies the minimal cost of transforming one distribution into another and effectively captures both location and shape differences, but it again relies on full distributions (Peyré & Cuturi, 2019). The MQC metric addresses a complementary objective: it evaluates the stability of probabilistic forecasts directly at the quantile level, without the need of full predictive densities. Since it is based solely on a finite set of quantiles rather than the full predictive distribution, MQC does not capture all aspects of the distributional change, potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles. However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest. Additionally, given that it is based on the pinball loss, the MQC is inherently asymmetric, reflecting the fact that quantile forecasts are directional. For example, an unstable 90% quantile forecast may have stronger implications on inventory or risk management than a similarly unstable 60% quantile forecast. This implies that MQC emphasizes directional changes when measuring stability, in contrast to symmetric distances (like $W_{2}$), which treat deviations identically regardless of the quantile’s position. Importantly, MQC is both model-agnostic and distribution-free, meaning that it can be applied consistently across different forecasting approaches without requiring strong assumptions about the underlying predictive distributions. Moreover, MQC is computationally lightweight compared to the Wasserstein distance or the KL divergence, which makes it suitable for large-scale applications involving thousands of time series.
We considered a total of 14 different quantiles of probability levels $\alpha\in$ {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we analyzed the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals. The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, essential to determine the safety stock levels in retail demand forecasting (Barrow & Kourentzes, 2016). These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution. To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of time series used 333 We used a validation set four times larger than the forecast horizon for the M5 dataset and twice the forecast horizon for the VN1 dataset. This limited the number of time series from 30,490 to 28,298 for the M5 dataset. .
We also normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining periods. To statistically validate our findings, we applied the Friedman-Nemenyi test for multiple comparisons (Demšar, 2006).
3.4.2 Performance metrics
Evaluating point forecast accuracy in time series is a debated topic, with no consensus on the best metric (Hewamalage et al., 2023). In our context of SKU-level demand forecasting, characterized by intermittent data, metrics based on absolute or percentage errors are suboptimal (Kolassa, 2020), and due to varying scales across series, a scaled accuracy metric has to be preferred. Hence, we adopted the Root Mean Squared Scaled Error (RMSSE) (Hyndman & Koehler, 2006),
which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (with $s=1$) (Makridakis et al., 2022b). Lower values indicate better performance.
These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016). Moreover, the SMQL was the official metric of the M5 Uncertainty competition (weighted and with $s=1$) (Makridakis et al., 2022c).
@@ 4 Results and discussion
Table 2: Overall forecasting performance and stability for each method across datasets based on the baseline retraining scenario, $r=7$ for M5 and $r=1$ for VN1. Minimum values per column are highlighted in bold.
Method M5 VN1
RMSSE SMQL sMAPC SMQC RMSSE SMQL sMAPC SMQC
LR 0.777 0.267 0.188 0.110 6.549 4.286 0.085 1.560
RF 1.868 1.220 0.092 0.729
XGBoost 0.755 0.245 0.075 0.100 1.890 1.503 0.146 0.629
LGBM 0.771 0.246 0.054 0.118 3.542 2.318 0.121 1.484
CatBoost 0.947 0.277 0.118 0.156 5.762 4.233 0.320 2.276
MLP 0.821 0.263 0.500 0.083 1.543 0.806 0.458 0.219
LSTM 1.913 0.892 0.558 0.242
TCN 0.865 0.270 0.499 0.083 1.913 0.891 0.600 0.241
NBEATSx 0.815 0.262 0.547 0.084 1.698 0.991 0.596 0.765
NHITS 0.828 0.266 0.566 0.087 1.699 0.992 0.605 0.755
Ens2A 0.757 0.243 0.059 0.108 1.472 0.826 0.503 0.441
Ens3A 0.757 0.248 0.089 0.107 1.517 0.867 0.514 0.539
Ens4A 0.758 0.240 0.095 0.099 1.524 0.926 0.167 0.545
Ens5A 0.763 0.241 0.099 0.095 1.544 1.013 0.164 0.540
Table 2 summarizes the performance and stability of all forecasting methods evaluated in this study, across both the M5 and VN1 datasets. For each method, the table reports four key metrics per dataset: RMSSE (Root Mean Squared Scaled Error) to evaluate point forecast accuracy, SMQL (Scaled Multi-Quantile Loss) to assess probabilistic forecast performance, sMAPC (symmetric Mean Absolute Percentage Change) to compare the stability in point forecast settings, and SMQC (Scaled Multi-Quantile Change) to understand the stability in probabilistic terms. Overall, the models evaluated achieved better stability results on the M5 dataset than on VN1. Several factors may account for this difference, such as the larger dataset size and a higher frequency of the M5 time series, the availability of rich external regressors (e.g., promotions, special events), and the existence of well-established benchmark hyperparameter settings, many of which were unavailable for VN1 at the time of the analysis. Moreover, in terms of point forecasting, machine learning models are more stable than deep learning methods on both datasets, implying that the former are less sensitive to frequent updates. On the contrary, in terms of probabilistic forecasting, deep learning architectures show less instability, meaning that these approaches can be a better option when retraining is performed frequently.
In general, ensembling seems to be a good solution to produce stable predictions when compared to base models. However, larger ensembles do not necessarily imply more stability, and it is not clear whether model diversity may have an impact on ensembles stability of predictions.
Figure 1 shows the point forecast stability of each model along the different retraining scenarios for the M5 and the VN1 datasets. To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M5 and $r=1$ for VN1. As expected, the sMAPC profiles are non-increasing functions of the retraining scenario. The stability remains practically the same across update frequencies or even improves as the retraining period increases. Indeed, moving from a high to a low retraining scenario, the stability of most global models improves compared to the baseline 444 Some models, like LR and MLP, show slightly decreasing stability (higher sMAPC) with an increasing retraining period. This can be due to many factors (e.g. sensitivity to small data changes, lack of robustness to stale parameters), but further investigation is necessary. . In particular, for some models (i.e. CatBoost, LSTM, and TCN), the instability reduction is strong and consistent, reaching very low levels. These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models. This can be explained by the fact that, when a forecasting origin is updated and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update. Therefore, less retraining is often synonymous with more stability.
@@
Figure 1: sMAPC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figures 2, and 3 statistically confirms the above results for the M5 and VN1 datasets. It is clear that most periodic retraining scenarios are statistically different to the continuous retraining strategy in terms of point forecast stability at the 5% level.
Figure 2: M5 Friedman-Nemenyi test results based on sMAPC.
Figure 3: VN1 Friedman-Nemenyi test results based on sMAPC.
In a similar fashion, Figure 4 summarizes the relative stability in a probabilistic forecasting setting. In this context, we observe that, for the M5 dataset, the stability (as measured by Scaled Multi Quantile Change) paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models 555 There are a few strange spikes in the SMQC profiles of some models. These can be due to different factors (e.g. inappropriate hyperparameter values), but further investigation is necessary. . For the VN1 dataset, instead, we observe some differences among the models. Some of them show constant improvements in stability as the retraining scenario increases (i.e. NBEATS, NHITS, and CatBoost), while most of them do not exhibit any significant enhancements or deteriorations in the retraining intervals. For both datasets, these results are confirmed by the statistical tests on SMQC over the different retraining periods (see Supplementary materials).
@@
Figure 4: SMQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Overall, the results from Figures 1 and 4, supported by the Friedman-Nemenyi tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts. When considered alongside the findings of Zanotti (2025b), this provides strong evidence against the practice of continuous retraining in global forecasting models. Lower retraining scenarios preserve both accuracy and stability, while offering substantial savings in computational resources.
Figure 5 shows the evolution of point forecast stability, across different retraining scenarios for the various ensemble configurations. The results on the M5 dataset indicate that reducing retraining scenario may negatively affect forecast stability. However, this impact is below 1% for most retraining scenarios. Indeed, statistical tests (in Supplementary material) on the significance of these differences show that most retraining scenarios are not statistically different from each other, with periodic retraining strategies being at least as good as continuous retraining. The results on the VN1, instead, indicate that there exists a clear benefit in lowering the retraining scenario in terms of forecast stability. Notably, on the M5 dataset the smallest ensemble (Ens2A) consistently achieves the lowest sMAPC values, since it is obtained by combining XGBoost and LGBM, that where both the most accurate and the most stable global forecasting models on that set of data. On the contrary, larger ensembles (such as Ens4A and Ens5A) stability profiles tend to outperform smaller ones in the VN1 dataset, suggesting that when base models are less accurate, bigger ensembles may be necessary to improve forecast stability.
Figure 6 further supports these observations in the probabilistic domain. First of all, probabilistic forecasts of ensembles are generally more stable than those of the base models across retraining frequencies, especially in M5, where SMQC variations are within 1 percentage point. In the VN1 dataset, the stability improvement is evident as the models are retrained less frequently. Then, from a probabilistic perspective, all the different ensembles perform similarly in terms of stability of the predictions, without showing any relevant benefit of having larger ensembles with respect to smaller ones. Most importantly, the effect of ensembling on the profiles of stability, as measured by sMAPC or SMQC, is that it greatly reduces the variations across the different retraining scenarios, reducing both positive and negative stability changes.
@@
Figure 5: sMAPC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figure 6: SMQC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figures 7 and 8 illustrate the fundamental trade-off between forecast accuracy and stability, comparing RMSSE with sMAPC and SMQL with SMQC, respectively. These plots reveal that high accuracy does not always guarantee high stability. In particular, several models achieve low RMSSE or SMQL scores but exhibit high instability, especially in VN1, highlighting their sensitivity to changes in the training data and their volatility across forecast origins. Nevertheless, it is the ensemble models that most consistently approach the optimality of stability and accuracy. In both the point and probabilistic domains, even very simple ensembles manage to control variance and mitigate instability without sacrificing much, if any, predictive performance. The overall picture that emerges from these figures reinforces the idea ensembling offers a practical and effective strategy to counteract the forecast instability.
Figure 7: The accuracy-stability trade-off. RMSSE vs sMAPC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
Figure 8: The accuracy-stability trade-off. SMQL vs SMQC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
This set of results has clear practical implications: in operational forecasting systems, especially in domains such as retail and supply chain management, the use of ensembles combined with lower retraining scenarios may offer a robust forecasting process, balancing performance with stability.

### v3

@@ 3.3 Evaluation strategy
Out-of-sample evaluation is a cornerstone of time series forecasting, providing a way to test a model’s generalization ability beyond the training period, an essential step given that future data may diverge from historical patterns due to structural breaks, level shifts, or unexpected shocks (Tashman, 2000). Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir and Benítez, 2012). It respects the chronological nature of time series data while enabling repeated assessments across multiple forecast origins. In this framework, the time series is split into a training set and a test set, where the model is trained on the former and evaluated on the latter. Forecasts are generated for a defined horizon $h$, and at each iteration, the forecast origin is shifted forward by a specified step size, typically one, to simulate real-time forecasting. The model is then retrained on the updated training data, using either a fixed-length or expanding window. Evaluation metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
Compared to fixed origin evaluations, rolling origin offers a more nuanced view of a model’s robustness by exposing it to a range of temporal conditions (including seasonal patterns, trends, and potential anomalies) that may not be captured in a single evaluation window (Bergmeir and Benítez, 2012). This iterative approach reduces the risk of overfitting to one specific train-test split, and is especially beneficial in operational contexts like retail, energy, logistics, or finance, where forecasts are regularly updated in response to incoming data. It also aligns well with real-world decision-making processes that require forecasts to adapt over time.
In practice, the design of a rolling origin evaluation depends on several parameters: the size of the test set, the length of the forecast horizon, the step size, and the windowing strategy (fixed vs expanding). The expanding window strategy is particularly suited for short time series, as it maximizes the amount of historical data available at each iteration. This approach is widely used in applied settings (Petropoulos and et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series. The Table 1 outlines the key parameters used in our experiments, including a test set spanning at least one full year, to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency. A step size of one is also relevant in the context of stability evaluation, since multiple forecasts for the same target are required to compute stability metrics.
@@
Furthermore, to investigate the impact of retraining on forecasting stability, we examined various retraining scenarios, or retraining windows. As defined in Zanotti (2025), each scenario $r$ indicates the number of new observations after which the model is retrained. These retraining scenarios are tailored to the data frequency, daily for M4 and M5, and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles. For instance, in the daily case, $r=7$ reflects weekly updates. Across all scenarios, we trained global models on each dataset, updating the model either completely at each retraining step or not at all. We excluded hyperparameter tuning due to its high cost and marginal expected benefit. We treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
@@ 3.4 Evaluation metrics
3.4.1 Performance metrics
Evaluating point forecast accuracy in time series is a debated topic, with no consensus on the best metric (Hewamalage et al., 2023). In our context, characterized by intermittent data, metrics based on absolute or percentage errors are suboptimal (Kolassa, 2020), and due to varying scales across series, a scaled accuracy metric should be preferred. Hence, we adopted the Root Mean Squared Scaled Error (RMSSE) (Hyndman and Koehler, 2006),
which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (with $s=1$) (Makridakis et al., 2022c). Lower values indicate better performance.
These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016). Moreover, the SMQL was the official metric of the M5 Uncertainty competition (weighted, and with $s=1$) (Makridakis et al., 2022a).
We considered a total of 14 different quantiles of probability levels $\alpha\in$ {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we analyzed the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals. The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails. These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution. To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of series used  [^2] We used a validation set four times larger than the forecast horizon for the M4 and M5 datasets, and twice the forecast horizon for the VN1 dataset. This limited the number of time series from 4,227 to 2,557 for the M4 dataset, and from 30,490 to 28,298 for the M5 dataset. .
We also normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining periods. To statistically validate our findings, we applied the Friedman-Nemenyi test (Demšar, 2006).
@@ 3.4.2 Stability metrics
Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and n1n-1 respectively. The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values. Lower values imply more stable predictions.
Beyond point forecasts, our study also emphasizes the probabilistic stability of the models under different retraining strategies. Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts. Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying retraining scenarios  [^3] One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, or to use inherently probabilistic models. These are also valid ways of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022a; Lainder and Wolfinger, 2022). However, comparing different probabilistic modeling frameworks is out of the scope of our analysis, and it may be an area of future research. . Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic quantile forecasts, we proposed a new measure, the Multi-Quantile Change (MQC), to comprehensively assess the stability of the probabilistic predictions. We defined the Quantile Change and the Multi-Quantile Change as:
@@
where $\mathcal{Q}$ is the set of possible quantile levels $\alpha$. The Quantile Change (QC) metric is a measure of the change in forecasted quantiles, $\hat{q}_{t}^{(\alpha)}$, between two consecutive forecast origins. Given two forecast origins, $n$ and n1n-1, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$. The formula mirrors the Quantile Loss but replaces the actual observations $y_{t}$, with previous quantile forecasts $\hat{q}_{t,n-1}^{(\alpha)}$, thereby shifting the focus from forecast accuracy to forecast stability. The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{q}_{t,n}^{(\alpha)}\approx\hat{q}_{t,n-1}^{(\alpha)}$), then QC will be small. Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of stability in its probabilistic output. By aggregating this across a set of quantile levels $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability, providing a comprehensive metric for probabilistic stability.
@@
To further contextualize the proposed metric, it is important to compare it with existing distributional distance metrics such as the Kullback–Leibler ($KL$) divergence and the Wasserstein distance ($W_{2}$). KL divergence measures the relative entropy between two probability distributions and is particularly sensitive to differences in the tails (Kullback and Leibler, 1951). However, it requires access to full predictive densities, which can limit its use in forecasting applications where outputs are often expressed only as quantiles. Wasserstein distance ($W_{2}$), by contrast, quantifies the minimal cost of transforming one distribution into another and effectively captures both location and shape differences, but it usually relies on full distributions (although it can be approximated through discretization) (Peyré and Cuturi, 2019). The MQC metric addresses a complementary objective: it evaluates the stability of probabilistic forecasts directly at the quantile level, without the need for full predictive densities. Since it is based solely on a finite set of quantiles rather than the full predictive distribution, MQC does not capture all aspects of the distributional change, potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles. However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest. Importantly, MQC is both model-agnostic and distribution-free, meaning that it can be applied consistently across different forecasting approaches without requiring strong assumptions about the underlying predictive distributions. Moreover, MQC is computationally lightweight compared to the Wasserstein distance or the KL divergence, which makes it suitable for large-scale applications involving thousands of time series.
@@ 4 Results and discussion
Figure 1: RMSSE and sMAPC results for each global method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M4 and M5 datasets, and $r=1$ for the VN1 dataset.
Figure 1 shows the point forecast accuracy (on the left) and the point forecast stability (on the right) of each model  [^4] For the long computing time required, the Random Forest model has not been trained on the M5 dataset, while the LSTM model has not been adopted on both the M4 and M5 datasets. Moreover, the CatBoost method has not been used on the M4 dataset since too few categorical variables are available for this dataset. along the different retraining scenarios for the M4, M5, and VN1 datasets. To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M4 and M5, and $r=1$ for VN1. In terms of forecast stability, the models’ behavior is similar across the different datasets. As expected, the sMAPC profiles are non-increasing functions of the retraining scenario. The stability remains practically the same, or even improves, as the retraining period increases. Indeed, moving from a high to a low retraining scenario, the stability of most global models improves compared to the baseline. In particular, for some models (i.e. CatBoost, LSTM, and TCN), the stability improvement is strong and consistent, reaching very low levels. These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models. This can be explained by the fact that, when a forecasting origin is updated, and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update. Therefore, less retraining is often synonymous with more stability. The statistical tests (see the Supplementary material) confirm the above results, showing that most periodic retraining scenarios are statistically different to the continuous retraining strategy in terms of point forecast stability at the 5% level. Similar considerations arise for the point forecast accuracy: the RMSSE profiles of the global models remain remarkably stable across retraining scenarios, sometimes displaying even modest improvements in accuracy when retrained less frequently (as is the case of deep learning models for the M4 dataset). In particular, under low periodic retraining schemes, the performance of most global models is statistically indistinguishable from the baseline configuration, and even under extreme settings (such as the no-retraining scenario) any degradation in accuracy remains below 5%. Overall, these findings show that global models can preserve, and in some cases even enhance, their forecasting accuracy over extended retraining intervals. Hence, frequent retraining may introduce unnecessary variability without delivering tangible gains in both predictive performance and forecast stability.
@@
Figure 2: SMQL and SMQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M4 and M5 datasets, and $r=1$ for the VN1 dataset.
In a similar fashion, Figure 2 summarizes the relative performance (on the left) and stability (on the right) from a probabilistic forecasting perspective. In this context, we observe that, for daily data, the stability paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models  [^5] There are a few spikes in the SMQC profiles of some models. These can be due to different factors (e.g., inappropriate hyperparameter values). Nevertheless, the profiles are stable overall. . For weekly data, instead, we observe some differences among the models. Some of them show constant improvements in stability as the retraining scenario increases, while others do not exhibit particular enhancements or deteriorations in most periodic retraining intervals. When shifting the focus from point to probabilistic forecasting, the impact of retraining frequency on accuracy becomes more pronounced. Unlike point forecasts, probabilistic accuracy is more sensitive to longer retraining intervals. In several cases, reducing the retraining frequency leads to a gradual deterioration in probabilistic performance, indicating that the estimation of predictive uncertainty is more affected by outdated model parameters than point predictions. Nevertheless, these effects remain limited for low periodic retraining scenarios, where differences in probabilistic accuracy are typically negligible (below 2%). Importantly, this sensitivity is not universal: some models exhibit clear accuracy losses, while others maintain stable probabilistic performance or even improve. Overall, these findings suggest that accuracy uncertainty estimates benefit more from regular updates, yet do not strictly require very frequent retraining to remain reliable.
@@
Figure 3: RMSSE, sMAPC, SMQL, and SMQC results for ETS method on the M5 dataset for each retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$.
To further contextualize the stability findings for global models, we also evaluated two classical local forecasting approaches: automatic ARIMA and automatic ETS  [^6] For the long computing time required, the ARIMA model has not been trained on the M5 dataset. . Both local models exhibit highly consistent behaviors across datasets and retraining scenarios (see Supplementary material). As an illustrative example, Figure 3 reports the results for ETS on the M5 dataset. The accuracy profiles for both point and probabilistic forecasting, measured by RMSSE and SMQL respectively, are strictly increasing functions of the retraining period, indicating a systematic deterioration in performance as retraining becomes less frequent. Conversely, as expected, the stability profiles, captured by sMAPC and SMQC, are strictly decreasing, showing that less frequent retraining leads to progressively more stable forecasts. This monotonic trade-off highlights an intrinsic property of local models: stability improvements are usually obtained at the cost of accuracy. In contrast, global models display markedly different dynamics, often maintaining both accuracy and stability over a wide range of retraining periods, and, in some cases, even improving along one dimension without compromising the other. This comparison suggests that global forecasting models are inherently more robust to reduced retraining, benefiting from cross-series learning that mitigates the sensitivity to individual time series updates, a property that local models, by construction, cannot exploit.
Overall, the results from Figures 1 and 2, supported by the statistical tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts. These results indicate that the accuracy–stability trade-off in global forecasting models is substantially milder than in local approaches. When considered alongside the findings of Zanotti (2025) on energy consumption, this provides strong evidence against the practice of frequent retraining in global forecasting models. Lower retraining scenarios preserve both accuracy and stability, while offering substantial savings in computational resources and energy consumption of forecasting systems.
Figure 4: The optimal retraining frequency identified for each forecasting objective, and for each series in the datasets considering all the models.
Finally, to jointly assess how accuracy- and stability-driven objectives translate into different retraining policies, Figure 4 provides a distributional view of the optimal retraining frequency across time series, in both point and probabilistic forecasting settings. Overall, the results indicate that continuous retraining is not required to achieve either accurate or stable forecasts when using global models. Across both point and probabilistic objectives, the retraining frequency can be safely reduced in favor of a periodic strategy, with stability-based criteria generally favoring slightly less frequent updates, in some cases even no retraining. However, compared to point forecasting, probabilistic objectives tend to favor slightly more frequent retraining, suggesting that accurate and stable uncertainty estimation requires a more regular model re-estimation. In practical terms, for daily data, a retraining interval of approximately 21–30 days emerges as a reasonable compromise across objectives, while for weekly data, retraining can be postponed usually up to 8 weeks without compromising performance. These retraining schedules strike a balance between the need to update models in response to potential distributional changes and the desire to limit unnecessary updates. Importantly, adopting such periodic retraining strategies can reduce the computational cost of forecasting by up to 75% Zanotti (2025), leading to more efficient and sustainable forecasting systems without sacrificing accuracy and also improving forecast stability.